# VideoCLIP-XL: Advancing Long Description Understanding for Video CLIP Models

### v1

@@ Abstract
Contrastive Language-Image Pre-training (CLIP) has been widely studied and applied in numerous applications. However, the emphasis on brief summary texts during pre-training prevent CLIP from understanding long descriptions. This issue is particularly acute regarding videos given that videos often contain abundant detailed contents. In this paper, we propose the VideoCLIP-XL (eXtra Length) model, which aims to unleash the long-description understanding capability of video CLIP models. Firstly, we establish an automatic data collection system and gather a large-scale VILD pre-training dataset111 https://huggingface.co/alibaba-pai/VILD. with VIdeo and Long-Description pairs. Then, we propose Text-similarity-guided Primary Component Matching (TPCM) to better learn the distribution of feature space while expanding the long description capability. We also introduce two new tasks namely Detail-aware Description Ranking (DDR) and Hallucination-aware Description Ranking (HDR) for further understanding improvement. Finally, we construct a Long Video Description Ranking (LVDR) benchmark222 https://huggingface.co/alibaba-pai/LVDR. for evaluating the long-description capability more comprehensively. Extensive experimental results on widely-used text-video retrieval benchmarks with both short and long descriptions and our LVDR benchmark can fully demonstrate the effectiveness of our method.333 https://huggingface.co/alibaba-pai/VideoCLIP-XL.
@@ 1 Introduction
The Contrastive Language-Image Pre-training (CLIP) model Radford et al. (2021) represents a pivotal development in the field of vision-language pre-training. It integrates text and image encoders to align these two modalities through contrastive learning. This methodology has been effectively applied in various applications, such as zero-shot classification Sun et al. (2023), text-image retrieval Luo et al. (2023), and text-to-image generation Rombach et al. (2022); Frans et al. (2022). However, one of the notable limitations of CLIP is its constrained capacity to process extensive textual descriptions, owing to its text encoder’s reliance on absolute positional embeddings. This limitation not only restricts the length of input text but also leads to poorer training outcomes for embeddings associated with higher token counts. Existing studies Zhang et al. (2024) have revealed a de facto effective token limit of just around 20.
Furthermore, the vanilla CLIP training procedure’s emphasis on brief summary texts compels the vision encoder to focus predominantly on the main features of the visual input, often overlooking smaller, yet potentially critical details. This issue is particularly acute in videos as compared to images, given that videos encapsulate a wealth of details across successive frames, along with additional information such as the sequence and flow of activities, camera movements, etc. In this context, existing video CLIP models Xu et al. (2021); Luo et al. (2022); Wang et al. (2023c) that employ vanilla CLIP training methodologies may struggle to accurately capture complex relationships and attributes, due to their reliance on a simplistic “bag of concepts” approach Tang et al. (2023b). To overcome these limitations, enhancing the model’s capability to comprehend long descriptions is crucial. Longer texts provide a rich tapestry of attributes and interconnections, offering a pathway to significantly improve the model’s performance and applicability in more complex scenarios.
@@ 3.1 VIdeo & Long-Description (VILD) Dataset
(c) Available Video & Long-Description Data. As previously mentioned, existing datasets pairing videos with long text descriptions are often limited by either the quantity or the domains/genres of videos. In this regard, we perform the Data Sampling operation over these datasets. Specifically, 57.8K video clips of gaming and city/scenic exploration scenarios in MiraData Ju et al. (2024) are all included in VLD. 50K long captions describing natural landscape are randomly sampled from Open-Sora-Dataset PKU-YuanGroup (2024). The Desc. Rewrite step is also involved at the end.
@@ 3.4 The New LVDR Benchmark
Hallucination is ubiquitous in contemporary LLMs and LMMs Liu et al. (2024a). Given a video, the video CLIP model with the ability to understand long texts should naturally possess the discernment to distinguish between correct and erroneous texts in long descriptions. To better evaluate such ability, we propose the Long Video Description Ranking (LVDR) benchmark dataset. We first randomly sample 2K video & long-description pairs from Shot2Story Han et al. (2023). Then, we perform a synthesis process similar to Fig. 3(a), iterating $p-1$ times and altering $q$ words during each iteration, and resulting in totally $p$ descriptions with gradually increasing degrees of hallucination. We denote such a subset as $p\times q$ and construct five subsets as $4\times 1$, $4\times 2$, $4\times 3$, $4\times 4$, $4\times 5$}. Each distinct subset is manually reviewed to avoid inappropriate replacement. Representative examples will be provided in Fig. 4. Based on our analysis, a better model needs to be able to correctly rank these descriptions in descending order of similarity given the video. Thus, we also design the evaluation criterion named ranking score (RS) which can be formulated as:
@@ A.1 Details of VILD Data Generation
During the data generation for pre-training, we leverage Qwen1.5-72B-Chat Bai et al. (2023) in LLM-based steps and LLaVA-v1.6-34B Liu et al. (2024b) in LMM-based steps. All prompts we used are listed as follows:
@@ A.4 Performance Comparison with More Models
As shown in Tab. 7, we involve more recent powerful and large cross-encoder models Li et al. (2023c); Wang et al. (2024) for comprehensive comparisons. Cross-encoder models, especially large multi-modal models (LMMs), typically add additional Transformer layers to model the deep interaction between vision and text representations. The model can generally boost the retrieval performance, while resulting in an unbearably slow retrieval speed when applied to the entire image/video collection since the cross-modal costs are required for each image/video sample whenever a new text query is given. In contrast, our VideoCLIP-XL which has the dual-encoder structure has obviously fewer parameters and retrieval time cost. Dual-encoder encodes the visual and textual inputs in a wholly decoupled manner. The vision representation is allowed to be pre-computed and re-used independent of the text queries. Such approaches can utilize fast approximate nearest neighbor (ANN) search Muja and Lowe (2009); Jegou et al. (2010); Johnson et al. (2019) at runtime to ensure high efficiency. For example, VideoCLIP-XL generally surpasses UMT-L Li et al. (2023c) on zero-shot text-video retrieval and has $\sim$4.14$\times$ faster retrieval speed on MSR-VTT without any bells and whistles, which can also indicate the effectiveness of our pre-training stage. It is also $\sim$8.69$\times$ faster than InternVideo2s2-1B. For fine-tuning, large cross-encoder models naturally surpass dual-encoder models owing to the cross-modal feature interaction. Yet, these models also suffer from the low inference speed issue, and hence can hardly be deployed in real-time applications.
@@
Table 7: R@1 scores of (a) zero-shot and (b) fine-tuned text-video retrieval on MSR-VTT Xu et al. (2016), LSMDC Rohrbach et al. (2015), DiDeMo Anne Hendricks et al. (2017), MSVD Chen and Dolan (2011), and ActivityNet Heilbron et al. (2015). [V:304M/T:124M/C:0M/47.53s] indicates that the vision encoder has 304M parameters, the text encoder has 124M parameters, the cross-encoder has 0M parameters, and this model needs 47.53s for text-video retrieval on MSR-VTT test set (1000 text-video pairs, tested on a single A100-80G GPU). The same goes for others.

### v2

@@ Abstract
Contrastive Language-Image Pre-training (CLIP) has been widely studied and applied in numerous applications. However, the emphasis on brief summary texts during pre-training prevents CLIP from understanding long descriptions. This issue is particularly acute regarding videos given that videos often contain abundant detailed contents. In this paper, we propose the VideoCLIP-XL (eXtra Length) model, which aims to unleash the long-description understanding capability of video CLIP models. Firstly, we establish an automatic data collection system and gather a large-scale VILD pre-training dataset111 https://huggingface.co/alibaba-pai/VILD. with VIdeo and Long-Description pairs. Then, we propose Text-similarity-guided Primary Component Matching (TPCM) to better learn the distribution of feature space while expanding the long description capability. We also introduce two new tasks namely Detail-aware Description Ranking (DDR) and Hallucination-aware Description Ranking (HDR) for further understanding improvement. Finally, we construct a Long Video Description Ranking (LVDR) benchmark222 https://huggingface.co/alibaba-pai/LVDR. for evaluating the long-description capability more comprehensively. Extensive experimental results on widely-used text-video retrieval benchmarks with both short and long descriptions and our LVDR benchmark can fully demonstrate the effectiveness of our method.333 https://huggingface.co/alibaba-pai/VideoCLIP-XL.
@@ 1 Introduction
The Contrastive Language-Image Pre-training (CLIP) model Radford et al. (2021) represents a pivotal development in the field of vision-language pre-training. It integrates text and image encoders to align these two modalities through contrastive learning. This methodology has been effectively applied in various applications, such as zero-shot classification Sun et al. (2023), text-image retrieval Luo et al. (2023), and text-to-image generation Rombach et al. (2022); Frans et al. (2022). However, one of the notable limitations of CLIP is its constrained capacity to process extensive textual descriptions, owing to its text encoder’s reliance on maximum positional embeddings with length 77. This limitation greatly restricts the length of input text, and existing studies Zhang et al. (2024) have also revealed a de facto effective token limit of just around 20.
Furthermore, the vanilla CLIP training procedure’s emphasis on brief summary texts compels the text/vision encoder to focus predominantly on the main features of the text/visual input, often overlooking smaller, yet potentially critical details. This issue is particularly acute in videos as compared to images, given that videos encapsulate a wealth of details across successive frames, along with additional information such as the sequence and flow of activities, camera movements, etc. In this context, existing video CLIP models Xu et al. (2021); Luo et al. (2022); Wang et al. (2023c) that employ vanilla CLIP training methodologies may struggle to accurately capture complex relationships and attributes, due to their reliance on a simplistic “bag of concepts” approach Tang et al. (2023b). To overcome these limitations, enhancing the model’s capability to comprehend long descriptions is crucial. Longer texts provide a rich tapestry of attributes and interconnections, offering a pathway to significantly improve the model’s performance and applicability in more complex scenarios.
@@ 3.1 VIdeo & Long-Description (VILD) Dataset
(c) Available Video & Long-Description Data. As previously mentioned, existing datasets pairing videos with long text descriptions are often limited by either the quantity or the domains/genres of videos. In this regard, we perform the Data Sampling operation over these datasets. Specifically, 57.8K video clips of gaming and city/scenic exploration scenarios in MiraData Ju et al. (2024) are all included in VILD. 50K long captions describing natural landscape are randomly sampled from Open-Sora-Dataset PKU-YuanGroup (2024). The Desc. Rewrite step is also involved at the end.
@@ 3.4 The New LVDR Benchmark
Hallucination is ubiquitous in contemporary LLMs and LMMs Liu et al. (2024a). Given a video, the video CLIP model with the ability to understand long texts should naturally possess the discernment to distinguish between correct and erroneous texts in long descriptions. To better evaluate such ability, we propose the Long Video Description Ranking (LVDR) benchmark. We first randomly sample 2K video & long-description pairs from Shot2Story Han et al. (2023). Then, we perform a synthesis process similar to Fig. 3(a), iterating $p-1$ times and altering $q$ words during each iteration, and resulting in totally $p$ descriptions with gradually increasing degrees of hallucination. We denote such a subset as $p\times q$ and construct five subsets as $4\times 1$, $4\times 2$, $4\times 3$, $4\times 4$, $4\times 5$}. Each distinct subset is manually reviewed to avoid inappropriate replacement. Representative examples are provided in Fig. 4. Based on our analysis, a better model needs to be able to correctly rank these descriptions in descending order of similarity given the video. Thus, we also design the evaluation criterion named ranking score (RS) which can be formulated as:
@@ A.1 Details of VILD Data Generation
During the data generation for VILD, we leverage Qwen1.5-72B-Chat Bai et al. (2023) in LLM-based steps and LLaVA-v1.6-34B Liu et al. (2024b) in LMM-based steps. All prompts we used are listed as follows:
@@ A.4 Performance Comparison with More Models
As shown in Tab. 7, we involve more recent powerful and large cross-encoder models Li et al. (2023c); Wang et al. (2024) for comprehensive comparisons. Cross-encoder models, especially large multi-modal models (LMMs), typically add additional Transformer layers to model the deep interaction between vision and text representations. The model can generally boost the retrieval performance, while resulting in an unbearably slow retrieval speed when applied to the entire image/video collection since the cross-modal costs are required for each image/video sample whenever a new text query is given. In contrast, our VideoCLIP-XL which has the dual-encoder structure has obviously fewer parameters and retrieval time cost. Dual-encoder encodes the visual and textual inputs in a wholly decoupled manner. The vision representation is allowed to be pre-computed and re-used independent of the text queries. Such approaches can utilize fast approximate nearest neighbor (ANN) search Muja and Lowe (2009); Jegou et al. (2010); Johnson et al. (2019) at runtime to ensure high efficiency. For example, VideoCLIP-XL generally surpasses UMT-L Li et al. (2023c) on zero-shot text-video retrieval and has $\sim$4.14$\times$ faster retrieval speed on MSR-VTT without any bells and whistles, which can also indicate the effectiveness of our pre-training stage. It is also $\sim$8.69$\times$ faster than InternVideo2s2-1B. For fine-tuning, large cross-encoder models naturally surpass dual-encoder models owing to the cross-modal feature interaction. Yet, these models still suffer from the low inference speed issue, and hence can hardly be deployed in real-time applications.
@@
Table 7: R@1 scores of (a) zero-shot and (b) fine-tuned text-video retrieval on MSR-VTT Xu et al. (2016), LSMDC Rohrbach et al. (2015), DiDeMo Anne Hendricks et al. (2017), MSVD Chen and Dolan (2011), and ActivityNet Heilbron et al. (2015). Due to the high overlap between the videos in ActivityNet and VideoInstruct100K Maaz et al. (2023), the latter is excluded from the pre-training data of our model tested on the former. [V:304M/T:124M/C:0M/47.53s] indicates that the vision encoder has 304M parameters, the text encoder has 124M parameters, the cross-encoder has 0M parameters, and this model needs 47.53s for text-video retrieval on MSR-VTT test set (1000 text-video pairs, tested on a single A100-80G GPU). The same goes for others.